If 1/2, −1 is the midpoint of segment QP and the coordinates of P are (− 1/2, 1), find the coordinates of Q.
Please help

Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



Use the midpoint formulas:

and



You are given and . Plug in the values and solve for and

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by stanbon(75887)

If (1/2, −1) is the midpoint of segment QP and the
coordinates of P are (− 1/2, 1), find the coordinates of Q.
-----------------
Let Q be (x,y)
----
(x+(-1/2))/2 =(1/2)
(2x-1)/4 = 1/2
2x-1 = 2
2x = 3
x = 3/2
--------------
(y+1)/2 = -1
y+1 = -2
y = -3
------
Q is (3/2,-3)
---------------
Cheers,
Stan H.